Fermi National Laboratory
 

Operation Summary for November 5, 2010 - November 8, 2010

 

The Integrated Luminosity for the period from 11/1/10 to 11/8/10 was 55.22 inverse picobarns. NuMI reported receiving 6.3E18 protons on target during this same period.

Friday Nov 5

The day shift began with Operations monitoring the Tevatron (TeV) store 8243 (55.65E30) and the Recycler stash (218.5E10), with the Antiproton Source (Pbar) stacking (20.7E10), with Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and Meson FTBF T-1004 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8243 at 8:30 AM, put the TeV into HEP shot setup, and established store 8244 at 10:15 AM with an average initial luminosity (AIL) of 245E30.

During the evening shift, Operations reported some minimal downtime due to two Linac RF stations (LRF1 and LRF3).

Saturday Nov 6

The midnight shift began with Operations monitoring store 8244 (54.11E30) and the stash (321.1E10), with Pbar stacking (17.6E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8244 at 3 AM, put the TeV into shot setup, and established store 8245 at 4:15 AM with an AIL of 334E30. Operations reported that a 400 MeV area Linac corrector magnet power supply developed a problem; Operations called in Linac personnel to replace the supply.

During the day shift, Operations monitored the accelerator complex.

During the evening shift, Operations terminated store 8245 at 4 PM. Operators prepared D0 for an access. After the access, operators put the TeV through a dry squeeze and then conducted orbit checks. Operations put the TeV into shot setup at 7:50 PM. Operations reported some jitter during Booster cogging. Operations called in Mechanical Support personnel to investigate a TeV sector A0 nitrogen tank pressure issue. Operations established store 8248 at 9 PM with an AIL of 344E30.

Sunday Nov 7

The midnight shift began with Operations monitoring store 8248 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, operators investigated NuMI beam position problems and Pbar Accumulator stacktail controls issues.

During the day shift, Operations terminated store 8248 at 10:15 AM, put the TeV into shot setup, and established store 8249 at 11:30 AM with an AIL of 325E30.

During the evening shift, Operations contacted NuMI personnel regarding several Lambertson (LAM61) trips.

Monday Nov 8

The midnight shift began with Operations monitoring store 8249 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8249 at 1:45 AM, put the TeV into shot setup, and established store 8250 at 3:15 AM with an AIL of 304E30. Operations reported some minimal downtime due to Linac and Booster RF.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and MI personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/08/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory